As a financial manager for WillPower, Inc, you have the following information:
a) The company follows a residual dividend policy;
b) The total capital budget for next year is likely to be $8,000,000;
c) The forecasted level of earnings next year is $8,000,000;
d) The target or optimal capital structure is a debt ratio of 40%;

Please answer the following questions:

a. What will be the amount funded with equity for the project ? (Keep the answer to a whole number. Example of answer format: $1,000,000)

b. Compute the amount of the dividend .   (Keep the answer to a whole number. Example of answer format: $1,000,000)

c. Compute the dividend pay-out ratio .     (Keep the answer to two decimals. Example of the answer format: 55.55%)
Answer #1

a. What will be the amount funded with equity for the project ?

Equity ==8000,000*.60=$4,800,000

Compute the amount of the dividend

Net income less amount funded with equity

dividend =$8,000,000-4,800,000=$3,200,000

Compute the dividend pay-out ratio

Dividend Payout ratio =Dividend/Earnings =3,200,000/8,000,000=40%